Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
menzione - 14:58 Thursday 18 July 2019 (46418) Print this report
Operator Report - Morning shift

05:00 UTC ITF in Science Mode with BNS Range 45 Mpc.
ITF unlocked for unknown reasons (plot1) at 05:45 UTC. It relocked at first attempt. Science Mode set again at 06:11 UTC. BNS Range ~45 Mpc.

Sub-system reports

Other
GRB
08:12 UTC [gracedb] Fermi event created: E344519. Alert reset at 08:13 UTC.

SQZ
At 11:16 UTC, SQZ unlocked. It re-ingaged it automatically. Science Mode set again at 11:28 UTC.

Images attached to this report
Comments to this report:
casanueva - 16:10 Thursday 18 July 2019 (46423) Print this report

The unlock reported by the operator is of the same kind as the ones from yesterday, an oscillation at 9.5 Hz that causes saturation in the B1 photodiode.

Images attached to this comment
ruggi - 18:14 Thursday 18 July 2019 (46424) Print this report

It is an oscillation of WE payload roll resonance. There is a damper which is supposed to suppress it, and in fact most of the time it seems to work, but sometimes for unknown reasons the oscillation grows; it looks like a loop close to the instability. It can stay excited for a while, then go down again and stay quiet for a long time. At the beginning of the run the WE roll damper was the only one not used; the 9.7 Hz peak was well visible in DARM, with some extra noise above 10 Hz. Then it was engaged after a small tuning, obtaining a certain improvement.

If now the occasional malfunctioning determines systematic unlocks, maybe it is better to go back and disengage the damper. Or maybe spend half an hour of commissioning time and try to tune better the loop.

ruggi - 21:04 Thursday 18 July 2019 (46426) Print this report

The attached plot shows the bad behavior at 9.7 Hz, happening from time to time. It was not so frequent, some time ago.

Images attached to this comment
ruggi - 12:43 Saturday 20 July 2019 (46439) Print this report

Considering that the latest locklosses belong to the 9.7 Hz cathegoty, the WE roll damper has been disengaged during the lock acquisition. Hopefully we will see the benefit soon.

ruggi - 16:25 Saturday 20 July 2019 (46441) Print this report

Fig 1 - the spectrogram shows the behavior of DARM at 9.7 Hz before and after the disengagement of the damper: now it is stably yellow, no more very unstable from green to red.

Fig 2 - the rms at 9.7 Hz of DARM and DIFFpTX is computed ad gives the same information.

Fig 3 - Hrec now is worsen at 10 Hz. No big effect on the quality of the data: it is much better to avoid locklosses. But we could try to tune the damper and have the good performance without the undesirable side effect.

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×